# Robust methods for Darcy–Forchheimer equationsDas, Hutridurga, Pani, Ruiz-Baier

### v1

@@ 1 Introduction
For Darcy and Forchheimer-type equations, which model creeping and inertial flows, respectively, the associated saddle-point systems exhibit strong sensitivity to variations in physical parameters such as permeability and the Forchheimer coefficient. Robust iterative solvers require preconditioners that maintain spectral equivalence independently of these parameters. For this, we follow the framework of operator preconditioning from [28] (see, also the earlier works [25, 2]), which suggests linking the preconditioning of algebraic systems with their infinite-dimensional operator counterpart. This allows to construct preconditioners by approximating the Riesz maps of the continuous operators in appropriate function space norms (or duality maps in the case of Banach spaces). For the mixed formulation of the Darcy problem, this approach yields block-diagonal preconditioners involving sums of spaces of pressure, as developed in [5] (and also extended for Biot equations and other families of perturbed saddle-point problems in, e.g., [8, 10, 23]). However, these works do not address in details the error analysis, which is also expected to maintain parameter robustness.
@@
The remainder of the paper is organised in the following manner. Section 2 recalls the necessary notations, the boundary value problem and the derivation of the weak formulation using parameter-weighted norms. Section 3 addresses the well-posedness of the weak formulation robustly with respect to model parameters, following the abstract theory of Minty–Browder monotone operators. In Section 4, we discuss the robust solvability of the discrete problem. In addition, we present quasi-optimal error analysis and derive precise convergence rates for conforming discretisations of the nonlinear model in Section 4.2. Next, in Section 5 we discuss how a simple variable operator preconditioning can be used in this context of Banach spaces, and we conclude in Section 6 with a set of numerical tests illustrating the properties of the proposed schemes.
@@ 2 Model problem
Let $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$, denote a bounded domain with Lipschitz-continuous boundary $\partial\Omega$, on which its outward unit normal is denoted by $\boldsymbol{n}$. For $s\geq 0$ and $r\in[1,+\infty]$, we denote by $\mathrm{L}^{r}(\Omega)$ and $\mathrm{W}^{s,r}(\Omega)$ the usual Lebesgue and Sobolev spaces endowed with the norms $\|\bullet\|_{0,r,\Omega}$ and $\|\bullet\|_{s,r,\Omega}$, respectively. Note that $\mathrm{W}^{0,r}(\Omega)=\mathrm{L}^{r}(\Omega)$. If $r=2$, we write $\mathrm{H}^{s}(\Omega)$ in place of $\mathrm{W}^{s,2}(\Omega)$, and denote the corresponding norm by $\|\bullet\|_{s,\Omega}$. We denote by $\mathbf{H}$ the vectorial counterpart of a generic scalar functional space $\mathrm{H}$. The $\mathrm{L}^{2}(\Omega)$ inner product for scalar, vector, or tensor valued functions is denoted by $(\bullet,\bullet)_{0,\Omega}$. We also recall the Hilbert space $\mathbf{H}(\mathop{\mathrm{div}}\nolimits,\Omega)=\{\boldsymbol{v}\in\mathbf{L}^{2}(\Omega):\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\in\mathrm{L}^{2}(\Omega)\}$ endowed with the norm $\|\boldsymbol{v}\|_{\mathop{\mathrm{div}}\nolimits,\Omega}^{2}=\|\boldsymbol{v}\|^{2}_{0,\Omega}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|^{2}_{0,\Omega}$. Moreover, we use for $r,s\in[1,+\infty)$, the Banach space
$\mathbf{H}^{r}(\mathop{\mathrm{div}}\nolimits_{s},\Omega)=\{\boldsymbol{v}\in\mathbf{L}^{r}(\Omega):\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\in\mathrm{L}^{s}(\Omega)\},$
endowed with the norm $\|\boldsymbol{v}\|_{r,\mathop{\mathrm{div}}\nolimits_{s},\Omega}^{2}=\|\boldsymbol{v}\|^{r}_{0,r,\Omega}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|^{s}_{0,s,\Omega}$. Often, we use a subscript $\Gamma_{i}$ on the functional space to denote that the (appropriate) traces vanish on the part of the boundary $\Gamma_{i}\subset\partial\Omega$.
@@
$\|r\|^{2}_{X\cap Y\cap Z}=\|r\|^{2}_{X}+\|r\|^{2}_{Y}+\|r\|^{2}_{Z},\quad\|z\|^{2}_{X+Y+Z}=\inf_{\begin{subarray}{c}r=u+v+w,\\ u\in X,v\in Y,w\in Z\end{subarray}}\,\bigl[\|u\|^{2}_{X}+\|v\|^{2}_{Y}+\|w\|^{2}_{Z}\bigr].$
@@
$a(\boldsymbol{u},\boldsymbol{v}):=\int_{\Omega}\kappa^{-1}\boldsymbol{u}\cdot\boldsymbol{v},\quad b(\boldsymbol{v},q):=\int_{\Omega}q\,\mathop{\mathrm{div}}\nolimits\boldsymbol{v},\quad F(\boldsymbol{v}):=\int_{\Omega}\boldsymbol{f}\cdot\boldsymbol{v},$
$G(q):=\int_{\Omega}g\,q,\quad c(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v}):=\int_{\Omega}\mathtt{F}|\boldsymbol{w}|^{r-2}\boldsymbol{u}\cdot\boldsymbol{v}.$
@@ 3 Continuous robust solvability
First, we recall an abstract setting and its solvability using the Minty–Browder framework, which we shall use it for the well-posedness of (2.3). A proof can be found in, e.g., [11, Theorem 3.1] (see also [12, Theorem 3.1]).
@@
(3.1) $\|(u,p)\|_{X\times Y}\leq C\bigl(\mathcal{M}(f,g)+\{\mathcal{M}(f,g)\}^{s-1}\bigr)$},
@@
We now remark that Theorem 3.1 readily yields the unique solvability of (2.3) if we specify the spaces $\mathbf{V}=\mathbf{H}^{3}_{\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)$ and $\mathrm{Q}=\mathrm{L}^{2}(\Omega)$ and this has been established in, e.g., [31]. However, the properties of the weak formulation (in particular, uniformly strong monotonicity on the kernel and inf-sup condition) are not necessarily uniform with respect to the model parameters, leading to lack of parameter robustness of the corresponding continuous dependence on data.
We adapt the operator preconditioning approach from [28] (and used for Darcy equations in [5]) and proceed first to show an auxiliary property of the divergence operator that will allow us to show a parameter-robust inf-sup condition.
@@
For the existence of ${\mathcal{S}_{c}}\in{\mathcal{L}}\bigl((\mathrm{W}_{\Gamma_{p}}^{1,q}(\Omega))^{\prime},{\mathbf{L}}^{r}(\Omega)\bigr)$ with $q\in\{2,\;\frac{3}{2}\}$, let $\phi\in\mathrm{W}^{1,r}_{\Gamma_{p}}(\Omega)$ for a given $g\in(\mathrm{W}_{\Gamma_{p}}^{1,q}(\Omega))^{\prime}$ with $\frac{1}{q}+\frac{1}{r}=1$ be a unique weak solution of the following elliptic problem with mixed boundary conditions (owing to, e.g., [30, Chapter 7])
@@
Then, it suffices to define ${\mathcal{S}}_{c}(g):=\nabla\phi$, where $\nabla\phi\in\mathbf{L}^{r}(\Omega)$. This, in turn, shows ${\mathcal{S}_{c}}\in{\mathcal{L}}\bigl((\mathrm{W}^{1,q}_{\Gamma_{p}}(\Omega))^{\prime},{\mathbf{L}}^{r}(\Omega)\bigr)$, for $q=2$ and $q=\frac{3}{2}$ corresponding to $r=2$ and $3,$ respectively.
@@
For a robust well-posedness of the Darcy–Forchheimer formulation, consider the following weighted spaces
$\mathbf{V}=\kappa^{-\frac{1}{2}}\mathbf{L}^{2}(\Omega)\cap\mathbf{H}_{\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)\cap\mathtt{F}^{\frac{1}{3}}\mathbf{L}^{3}(\Omega),\quad\mathrm{Q}=\mathrm{L}^{2}(\Omega)+\kappa^{\frac{1}{2}}\mathrm{H}^{1}_{\Gamma_{p}}(\Omega)+\mathtt{F}^{-\frac{1}{3}}\mathrm{W}^{1,\frac{3}{2}}_{\Gamma_{p}}(\Omega),$
endowed with the following norms
(3.6a) $\|\boldsymbol{v}\|_{\mathbf{V}}$ $:=\kappa^{-\frac{1}{2}}\|\boldsymbol{v}\|_{0,\Omega}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|_{0,\Omega}+\mathtt{F}^{\frac{1}{3}}\|\boldsymbol{v}\|_{0,3,\Omega},$
(3.6b) $\|q\|_{\mathrm{Q}}$ $:=\!\!\inf_{\begin{subarray}{c}q=q_{1}+q_{2}+q_{3},\\ (q_{1},q_{2},q_{3})\in\mathrm{L}^{2}(\Omega)\times\mathrm{H}^{1}_{\Gamma_{\!p}}\!(\Omega)\times\mathrm{W}^{1,\frac{3}{2}}_{\Gamma_{p}}(\Omega)\end{subarray}}\!\!\bigl[\|q_{1}\|_{0,\Omega}+\kappa^{\frac{1}{2}}\|q_{2}\|_{1,\Omega}+\mathtt{F}^{-\frac{1}{3}}\|q_{3}\|_{1,\frac{3}{2},\Omega}\bigr],$
respectively.
@@
Assume that $\boldsymbol{f}\in\mathbf{L}^{2}(\Omega),g\in\mathrm{L}^{2}(\Omega).$ Then, there exists a unique pair of solutions $(\boldsymbol{u},p)\in\mathbf{V}\times\mathrm{Q}$ to the problem (2.3). Furthermore, there is a positive constant $\tilde{C}$, independent of the model parameters $\kappa,\mathtt{F}$, such that
$\|(\boldsymbol{u},p)\|_{\mathbf{V}\times\mathrm{Q}}\leq\tilde{C}\max\bigl\{(\|\boldsymbol{f}\|_{0,\Omega}+\|g\|_{0,\Omega}+\|g\|^{2}_{0,\Omega}),(\|\boldsymbol{f}\|_{0,\Omega}+\|g\|_{0,\Omega}+\|g\|^{2}_{0,\Omega})^{2}\bigr\}.$
@@
In order to apply Theorem 3.1, we define the operators $A:\mathbf{V}\to\mathbf{V}^{\prime}$, $B:\mathbf{V}\to\mathrm{Q}^{\prime}$, and $C:\mathbf{V}\to\mathbf{V}^{\prime}$ as $\langle A\boldsymbol{u},\boldsymbol{w}\rangle:=a(\boldsymbol{u},\boldsymbol{w})$, $\langle B\boldsymbol{v},p\rangle:=b(\boldsymbol{v},p)$, and $\langle C(\boldsymbol{u}),\boldsymbol{w}\rangle:=c(\boldsymbol{u};\boldsymbol{u},\boldsymbol{w})$, respectively. Now, set the non-linear operator $A+C:\mathbf{V}\to\mathbf{V}^{\prime}$. We proceed to verify the assumptions of Theorem 3.1 identifying $X\equiv\mathbf{V}$, $Y\equiv\mathrm{Q}$, $\mathcal{A}\equiv A+C$, $\mathcal{B}\equiv B$, and $s=3$.
@@
$\langle\{A+C\}({\boldsymbol{u}}),\boldsymbol{w}\rangle-\langle\{A+C\}({\boldsymbol{v}}),\boldsymbol{w}\rangle=\int_{\Omega}\kappa^{-1}(\boldsymbol{u}-\boldsymbol{v})\cdot\boldsymbol{w}+\int_{\Omega}\mathtt{F}(|\boldsymbol{u}|\boldsymbol{u}-|\boldsymbol{v}|\boldsymbol{v})\cdot\boldsymbol{w}$
$\quad\leq\kappa^{-\frac{1}{2}}\|{\boldsymbol{u}}-{\boldsymbol{v}}\|_{0,\Omega}\kappa^{-\frac{1}{2}}\|\boldsymbol{w}\|_{0,\Omega}+\mathtt{F}^{\frac{1}{3}}\|\boldsymbol{w}\|_{0,3,\Omega}\mathtt{F}^{\frac{2}{3}}\||\boldsymbol{u}|\boldsymbol{u}-|\boldsymbol{v}|\boldsymbol{v}\|_{0,\frac{3}{2},\Omega}.$
@@
$\leq\kappa^{-\frac{1}{2}}\|{\boldsymbol{u}}-{\boldsymbol{v}}\|_{0,\Omega}\kappa^{-\frac{1}{2}}\|\boldsymbol{w}\|_{0,\Omega}+\mathtt{F}^{\frac{1}{3}}\|\boldsymbol{w}\|_{0,3,\Omega}\tilde{C}\mathtt{F}^{\frac{1}{3}}(\|\boldsymbol{u}\|_{0,3,\Omega}+\|\boldsymbol{v}\|_{0,3,\Omega})\mathtt{F}^{\frac{1}{3}}\|\boldsymbol{u}-\boldsymbol{v}\|_{0,3,\Omega}$
@@
$\quad=\int_{\Omega}\kappa^{-1}|\boldsymbol{u}-\boldsymbol{v}|^{2}+\mathtt{F}\int_{\Omega}\left((\boldsymbol{u}+\boldsymbol{w})|\boldsymbol{u}+\boldsymbol{w}|-(\boldsymbol{v}+\boldsymbol{w})|\boldsymbol{v}+\boldsymbol{w}|\right)\cdot(\boldsymbol{u}-\boldsymbol{v}).$
@@
$\mathtt{F}\int_{\Omega}\left((\boldsymbol{u}+\boldsymbol{w})|\boldsymbol{u}+\boldsymbol{w}|-(\boldsymbol{v}+\boldsymbol{w})|\boldsymbol{v}+\boldsymbol{w}|\right)\cdot(\boldsymbol{u}-\boldsymbol{v})\geq\tilde{C}\,\mathtt{F}\|\boldsymbol{u}-\boldsymbol{v}\|^{3}_{0,3,\Omega}$
@@
where $\beta=\|\mathcal{S}_{c}\|^{-1}_{\mathcal{L}(Q^{\prime},\mathbf{V})}$ is independent of the parameters $\kappa,\mathtt{F}$, and hence, the inf-sup condition is satisfied.
An appeal to the Theorem 3.1 implies the existence of a unique velocity-pressure pair of solutions to the problem (2.3).
@@
(3.9) $\|p\|_{\mathrm{Q}}\leq\frac{1}{\beta}\|B^{*}(p)\|_{\mathbf{V}}\leq\{\|{F}\|_{\mathbf{V}^{\prime}}+\|\{A+C\}(\boldsymbol{u})\|_{\mathbf{V}^{\prime}}\}.$
@@
[…]
Now, for the uniformly strong monotonicity of $A+C$ (cf. (3.8)), we choose $\boldsymbol{u}=\widetilde{\boldsymbol{u}},\boldsymbol{v}=\boldsymbol{0}$ and $\boldsymbol{w}=-\hat{\boldsymbol{u}}$, and then for equation (2.3a), we set $\boldsymbol{v}=\widetilde{\boldsymbol{u}}$. Further, we use the local Lipschitz continuity of $A$, by setting $\boldsymbol{v}=\boldsymbol{0}$ in (3.7), to conclude that
$\alpha\|\widetilde{\boldsymbol{u}}\|^{2}_{\mathbf{V}}$ $\leq\langle\{A+C\}(\widetilde{\boldsymbol{u}}+\hat{\boldsymbol{u}})-\{A+C\}(\boldsymbol{0}+\hat{\boldsymbol{u}}),\widetilde{\boldsymbol{u}}\rangle$
$\leq\{\|{F}\|_{\mathbf{V}^{\prime}}+\|\{A+C\}(\hat{\boldsymbol{u}})\|_{\mathbf{V}^{\prime}}\}\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}}$
$\leq\{\|{F}\|_{\mathbf{V}^{\prime}}+\|\hat{\boldsymbol{u}}\|_{\mathbf{V}}+\|\hat{\boldsymbol{u}}\|_{\mathbf{V}}^{2}\}\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}}\leq\bigl\{\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{2}}\|G\|_{\mathrm{Q}^{\prime}}^{2}\bigr\}\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}},$
and hence,
(3.10) $\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}}\leq\frac{1}{\alpha}\bigl\{\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{2}}\|G\|_{\mathrm{Q}^{\prime}}^{2}\bigr\}.$
Finally, from (3.10), using the triangle inequality, we find that
(3.11) $\|\boldsymbol{u}\|_{\mathbf{V}}\leq\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}}+\|\hat{\boldsymbol{u}}\|_{\mathbf{V}}\leq\frac{1}{\alpha}\bigl\{\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{2}}\|G\|_{\mathrm{Q}^{\prime}}^{2}\bigr\}.$
On the other hand, (3.9) and (2.3a) lead to

### v2

@@ 1 Introduction
For Darcy and Forchheimer-type equations, which model creeping and inertial flows, respectively, the associated saddle-point systems exhibit strong sensitivity to variations in physical parameters such as permeability and the Forchheimer coefficient. Robust iterative solvers require preconditioners that maintain spectral equivalence independently of these parameters. For this, we follow the framework of operator preconditioning from [29] (see, also the earlier works [26, 2]), which suggests linking the preconditioning of algebraic systems with their infinite-dimensional operator counterpart. This allows to construct preconditioners by approximating the Riesz maps of the continuous operators in appropriate function space norms (or duality maps in the case of Banach spaces). For the mixed formulation of the Darcy problem, this approach yields block-diagonal preconditioners involving sums of spaces of pressure, as developed in [5] (and also extended for Biot equations and other families of perturbed saddle-point problems in, e.g., [9, 8, 11, 24]). However, these works do not address in details the error analysis, which is also expected to maintain parameter robustness.
@@
The paper is organised as follows. Section 2 recalls the necessary notations, the boundary value problem and the derivation of the weak formulation using parameter-weighted norms. Section 3 addresses the well-posedness of the weak formulation robustly with respect to model parameters, following the abstract theory of Minty–Browder monotone operators. In Section 4, we discuss the robust solvability of the discrete problem. In addition, we present quasi-optimal error analysis and derive precise convergence rates for conforming discretisations of the nonlinear model in Section 4.2. Next, in Section 5 we discuss how a simple variable operator preconditioning can be used in this context of Banach spaces, and we conclude in Section 6 with a set of numerical tests illustrating the properties of the proposed schemes.
@@ 2 Model problem
Let $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$, denote a bounded domain with Lipschitz-continuous boundary $\partial\Omega$, on which its outward unit normal is denoted by $\boldsymbol{n}$. For $s\geq 0$ and $t\in[1,+\infty]$, we denote by $\mathrm{L}^{t}(\Omega)$ and $\mathrm{W}^{s,t}(\Omega)$ the usual Lebesgue and Sobolev spaces endowed with the norms $\|\bullet\|_{0,t,\Omega}$ and $\|\bullet\|_{s,t,\Omega}$, respectively. Note that $\mathrm{W}^{0,t}(\Omega)=\mathrm{L}^{t}(\Omega)$. If $t=2$, we write $\mathrm{H}^{s}(\Omega)$ in place of $\mathrm{W}^{s,2}(\Omega)$, and denote the corresponding norm by $\|\bullet\|_{s,\Omega}$. We denote by $\mathbf{H}$ the vectorial counterpart of a generic scalar functional space $\mathrm{H}$. The $\mathrm{L}^{2}(\Omega)$ inner product for scalar, vector, or tensor valued functions is denoted by $(\bullet,\bullet)_{0,\Omega}$. We also recall the Hilbert space $\mathbf{H}(\mathop{\mathrm{div}}\nolimits,\Omega)=\{\boldsymbol{v}\in\mathbf{L}^{2}(\Omega):\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\in\mathrm{L}^{2}(\Omega)\}$ endowed with the norm $\|\boldsymbol{v}\|_{\mathop{\mathrm{div}}\nolimits,\Omega}^{2}=\|\boldsymbol{v}\|^{2}_{0,\Omega}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|^{2}_{0,\Omega}$. Moreover, we use for $t,s\in[1,+\infty)$, the Banach space
$\mathbf{H}^{t}(\mathop{\mathrm{div}}\nolimits_{s},\Omega)=\{\boldsymbol{v}\in\mathbf{L}^{t}(\Omega):\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\in\mathrm{L}^{s}(\Omega)\},$}
endowed with the norm $\|\boldsymbol{v}\|_{t,\mathop{\mathrm{div}}\nolimits_{s},\Omega}^{2}=\|\boldsymbol{v}\|^{t}_{0,t,\Omega}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|^{s}_{0,s,\Omega}$. Often, we use a subscript $\Gamma_{i}$ on the functional space to denote that the (appropriate) traces vanish on the part of the boundary $\Gamma_{i}\subset\partial\Omega$.
@@
$\|r\|^{2}_{X\cap Y\cap Z}=\|r\|^{2}_{X}+\|r\|^{2}_{Y}+\|r\|^{2}_{Z},\quad\|z\|^{2}_{X+Y+Z}=\inf_{\begin{subarray}{c}r=u+v+w,\\ u\in X,v\in Y,w\in Z\end{subarray}}\,\bigl[\|u\|^{2}_{X}+\|v\|^{2}_{Y}+\|w\|^{2}_{Z}\bigr],$
and we also use the convenient equivalent form
$\|r\|_{X\cap Y\cap Z}=\|r\|_{X}+\|r\|_{Y}+\|r\|_{Z},\quad\|z\|_{X+Y+Z}=\inf_{\begin{subarray}{c}r=u+v+w,\\ u\in X,v\in Y,w\in Z\end{subarray}}\,\bigl[\|u\|_{X}+\|v\|_{Y}+\|w\|_{Z}\bigr].$
@@
$a(\boldsymbol{u},\boldsymbol{v}):=\int_{\Omega}\kappa^{-1}\boldsymbol{u}\cdot\boldsymbol{v},\quad b(\boldsymbol{v},q):={-\int_{\Omega}q\,\mathop{\mathrm{div}}\nolimits\boldsymbol{v}},\quad F(\boldsymbol{v}):=\int_{\Omega}\boldsymbol{f}\cdot\boldsymbol{v},$
$G(q):={-\int_{\Omega}g\,q},\quad c(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v}):=\int_{\Omega}\mathtt{F}|\boldsymbol{w}|^{r-2}\boldsymbol{u}\cdot\boldsymbol{v}.$
@@ 3 Continuous robust solvability
First, we recall an abstract setting and its solvability using the Minty–Browder framework, which we use for the well-posedness of (2.3). A proof can be found in, e.g., [12, Theorem 3.1] (see also [13, Theorem 3.1]).
@@
(3.1) $\|(u,p)\|_{X\times Y}\leq C\bigl(\mathcal{M}(f,g)+{\bigl(\mathcal{M}(f,g)\bigr)}^{s-1}\bigr)$},
@@
Theorem 3.1 readily yields the unique solvability of (2.3) if we specify the spaces $\mathbf{V}=\mathbf{H}^{{r}}_{\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)$ and $\mathrm{Q}=\mathrm{L}^{2}(\Omega)$. This has been established in, e.g., [32] for the case $r=3$. However, the properties of the weak formulation (in particular, uniformly strong monotonicity on the kernel and inf-sup condition) are not necessarily uniform with respect to the model parameters, leading to lack of parameter robustness of the corresponding continuous dependence on data.
For a robust well-posedness of the Darcy–Forchheimer formulation, we consider the following weighted spaces
$\mathbf{V}:=\kappa^{-\frac{1}{2}}\mathbf{L}^{2}(\Omega)\cap\mathbf{H}_{\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)\cap{\mathtt{F}^{\frac{1}{r}}\mathbf{L}^{r}(\Omega)},\quad\mathrm{Q}:=\mathrm{L}^{2}(\Omega)+\kappa^{\frac{1}{2}}\mathrm{H}^{1}_{\Gamma_{p}}(\Omega)+{\mathtt{F}^{-\frac{1}{r}}\mathrm{W}^{1,r^{\prime}}_{\Gamma_{p}}(\Omega)},$
endowed with the following norms
(3.2a) $\|\boldsymbol{v}\|_{\mathbf{V}}$ $:=\kappa^{-\frac{1}{2}}\|\boldsymbol{v}\|_{0,\Omega}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|_{0,\Omega}+{\mathtt{F}^{\frac{1}{r}}\|\boldsymbol{v}\|_{0,r,\Omega}},$
(3.2b) $\|q\|_{\mathrm{Q}}$ $:=\!\!\inf_{\begin{subarray}{c}q=q_{1}+q_{2}+q_{3},\\ (q_{1},q_{2},q_{3})\in\mathrm{L}^{2}(\Omega)\times\mathrm{H}^{1}_{\Gamma_{\!p}}\!(\Omega)\times{\mathrm{W}^{1,r^{\prime}}_{\Gamma_{p}}(\Omega)}\end{subarray}}\!\!\bigl[\|q_{1}\|_{0,\Omega}+\kappa^{\frac{1}{2}}\|q_{2}\|_{1,\Omega}+{\mathtt{F}^{-\frac{1}{r}}|q_{3}|_{1,r^{\prime},\Omega}}\bigr],$
respectively. This choice becomes evident in the analysis to follow. We adapt the operator preconditioning approach from [29] (and used for Darcy equations in [5]) and proceed first to show an auxiliary property of the divergence operator that allows us to prove a parameter-robust inf-sup condition.
@@
Recall that $r\in[3,4]$, and let $r^{\prime}$ be such $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. For the existence of ${\mathcal{S}_{c}}\in{\mathcal{L}}\bigl((\mathrm{W}_{\Gamma_{p}}^{1,q}(\Omega))^{\prime},{\mathbf{L}}^{{q^{\prime}}}(\Omega)\bigr)$ with $q\in\{2,\;{r^{\prime}}\}$ and $\frac{1}{q}+\frac{1}{q^{\prime}}=1$, let $\phi\in\mathrm{W}^{1,{q^{\prime}}}_{\Gamma_{p}}(\Omega)$ for a given $g\in(\mathrm{W}_{\Gamma_{p}}^{1,q}(\Omega))^{\prime}$ be the unique weak solution of the following elliptic problem with mixed boundary conditions (owing to, e.g., [31, Chapter 7])
@@
Then, it suffices to define ${\mathcal{S}}_{c}(g):=\nabla\phi$, where $\nabla\phi\in\mathbf{L}^{{q^{\prime}}}(\Omega)$. This, in turn, shows ${\mathcal{S}_{c}}\in{\mathcal{L}}\bigl((\mathrm{W}^{1,q}_{\Gamma_{p}}(\Omega))^{\prime},{\mathbf{L}}^{{q^{\prime}}}(\Omega)\bigr)$, for $q=2$ and $q={r^{\prime}}$ corresponding to $q^{\prime}=2$ and $q^{\prime}=r$, respectively.
@@
Assume that $\boldsymbol{f}\in\mathbf{L}^{2}(\Omega),g\in\mathrm{L}^{2}(\Omega).$ Then, there exists a unique $(\boldsymbol{u},p)\in\mathbf{V}\times\mathrm{Q}$ solution to (2.3). Furthermore, there is a positive constant $\tilde{C}$, independent of the model parameters $\kappa,\mathtt{F}$, such that
$\|(\boldsymbol{u},p)\|_{\mathbf{V}\times\mathrm{Q}}\leq\tilde{C}\max\bigl\{\|\boldsymbol{f}\|_{0,\Omega}+\|g\|_{0,\Omega}+{\|g\|^{r-1}_{0,\Omega}},(\|\boldsymbol{f}\|_{0,\Omega}+\|g\|_{0,\Omega}+{\|g\|^{r-1}_{0,\Omega}})^{2}\bigr\}.$
@@
In order to apply Theorem 3.1, we define the operators $A:\mathbf{V}\to\mathbf{V}^{\prime}$, $B:\mathbf{V}\to\mathrm{Q}^{\prime}$, and $C:\mathbf{V}\to\mathbf{V}^{\prime}$ as $\langle A\boldsymbol{u},\boldsymbol{w}\rangle:=a(\boldsymbol{u},\boldsymbol{w})$, $\langle B\boldsymbol{v},p\rangle:=b(\boldsymbol{v},p)$, and $\langle C(\boldsymbol{u}),\boldsymbol{w}\rangle:=c(\boldsymbol{u};\boldsymbol{u},\boldsymbol{w})$, respectively. Now, set the non-linear operator $A+C:\mathbf{V}\to\mathbf{V}^{\prime}$. We proceed to verify the assumptions of Theorem 3.1 identifying $X\equiv\mathbf{V}$, $Y\equiv\mathrm{Q}$, $\mathcal{A}\equiv A+C$, $\mathcal{B}\equiv B$, and $s=r$.
@@
$=\int_{\Omega}\kappa^{-1}(\boldsymbol{u}-\boldsymbol{v})\cdot\boldsymbol{w}+\int_{\Omega}\mathtt{F}({|\boldsymbol{u}|^{r-2}}\boldsymbol{u}-{|\boldsymbol{v}|^{r-2}}\boldsymbol{v})\cdot\boldsymbol{w}$
$\quad\leq\kappa^{-\frac{1}{2}}\|{\boldsymbol{u}}-{\boldsymbol{v}}\|_{0,\Omega}\kappa^{-\frac{1}{2}}\|\boldsymbol{w}\|_{0,\Omega}+{\mathtt{F}^{\frac{1}{r}}\|\boldsymbol{w}\|_{0,r,\Omega}\mathtt{F}^{\frac{1}{r^{\prime}}}\||\boldsymbol{u}|^{r-2}\boldsymbol{u}-|\boldsymbol{v}|^{r-2}\boldsymbol{v}\|_{0,r^{\prime},\Omega}}.$
@@
$\leq\kappa^{-\frac{1}{2}}\|{\boldsymbol{u}}-{\boldsymbol{v}}\|_{0,\Omega}\kappa^{-\frac{1}{2}}\|\boldsymbol{w}\|_{0,\Omega}+{\mathtt{F}^{\frac{1}{r}}\|\boldsymbol{w}\|_{0,r,\Omega}\tilde{C}\mathtt{F}^{\frac{1}{r}}(\|\boldsymbol{u}\|_{0,r,\Omega}+\|\boldsymbol{v}\|_{0,r,\Omega})\mathtt{F}^{\frac{1}{r}}\|\boldsymbol{u}-\boldsymbol{v}\|_{0,r,\Omega}}$
@@
$\quad=\int_{\Omega}\kappa^{-1}|\boldsymbol{u}-\boldsymbol{v}|^{2}+\mathtt{F}\int_{\Omega}{\left((\boldsymbol{u}+\boldsymbol{w})|\boldsymbol{u}+\boldsymbol{w}|^{r-2}-(\boldsymbol{v}+\boldsymbol{w})|\boldsymbol{v}+\boldsymbol{w}|^{r-2}\right)}\cdot(\boldsymbol{u}-\boldsymbol{v}).$
@@
$\mathtt{F}\int_{\Omega}\left((\boldsymbol{u}+\boldsymbol{w})|\boldsymbol{u}+\boldsymbol{w}|^{r-2}-(\boldsymbol{v}+\boldsymbol{w})|\boldsymbol{v}+\boldsymbol{w}|^{r-2}\right)\cdot(\boldsymbol{u}-\boldsymbol{v})\geq\tilde{C}\,{\mathtt{F}\|\boldsymbol{u}-\boldsymbol{v}\|^{r}_{0,r,\Omega}.}$
@@
where $\beta=\|\mathcal{S}_{c}\|^{-1}_{\mathcal{L}(Q^{\prime},\mathbf{V})}$ is independent of the parameters $\kappa,\mathtt{F}$, and hence, the inf-sup condition is satisfied. Note also that the boundedness of $B$ follows directly from Hólder’s inequality
$|b(\boldsymbol{v},q)|=|b(\kappa^{-\frac{1}{2}}\boldsymbol{v},\kappa^{\frac{1}{2}}q)|\leq\kappa^{-\frac{1}{2}}\|\boldsymbol{v}\|_{0,\Omega}\kappa^{\frac{1}{2}}\|q\|_{0,\Omega}\leq\|\boldsymbol{v}\|_{\mathbf{V}}\|q\|_{\mathrm{Q}}.$
Therefore, an appeal to the Theorem 3.1 implies the existence of a unique velocity-pressure pair of solutions to the problem (2.3).
@@
(3.9) $\|p\|_{\mathrm{Q}}\leq\frac{1}{\beta}\|B^{*}(p)\|_{\mathbf{V}}\leq{\|{F}\|_{\mathbf{V}^{\prime}}+\|\{A+C\}(\boldsymbol{u})\|_{\mathbf{V}^{\prime}}}.$